OpenStudy (webo):

Solve −2x − 13 = 8x + 7.

OpenStudy (acespeedfighter):

nice profile pic

OpenStudy (kgrendel0324):

Thnx :D

OpenStudy (kgrendel0324):

get the variable on one side of the equation and the constants on the other.

alones (alones):

Subtract 8x from both sides. −2x−13−8x=8x+7−8x −10x−13=7 than add 13 to both sides. −10x−13+13=7+13 −10x=20

alones (alones):

After all ivide both sides by -10, and that should be ur answer, tell me what you'll get

OpenStudy (kgrendel0324):

Alone, I tried, he isnt there.